OpenStudy (anonymous):

Express as a single logarithm. Simplify, if possible. log5.4 - log0.054

OpenStudy (anonymous):

log (5.4) * log (0.054) = log (0.2916)

OpenStudy (phi):

log5.4 - log0.054 Do you know the rule log(a)-log(b)= log(a/b) ? if you do, then you know log5.4 - log0.054 = log(5.4/0.054)= log(1/100) log means log base 10. Can you finish?